OpenStudy (anonymous):

Write the standard form of the equation of the circle with the given characteristics: Center(-5,-6) radius=6... Pleaseeeeeeeeeeeeeeeeeeeeeee

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] with \(h=-5,k=-6,r=6\)

OpenStudy (anonymous):

I already replaced h, k, and r with -5, -6, and 6... I still don't get it right..... What else do I have to do after plugging in the numbers.... I have (x+5)^2-(y+6)^2= 36.... Help